# lemmasection

### v1

@@ Abstract
ECG foundation models are increasingly popular due to their adaptability across various tasks. However, their clinical applicability is often limited by performance gaps compared to task-specific models, even after pre-training on large ECG datasets and fine-tuning on target data. This limitation is likely due to the lack of an effective post-training strategy. In this paper, we propose a simple yet effective post-training approach to enhance ECGFounder, a state-of-the-art ECG foundation model pre-trained on over 7 million ECG recordings. Experiments on the PTB-XL benchmark show that our approach improves the baseline fine-tuning strategy by 1.2%–3.3% in macro AUROC and 5.3%–20.9% in macro AUPRC. Additionally, our method outperforms several recent state-of-the-art approaches, including task-specific and advanced architectures. Further evaluation reveals that our method is more stable and sample-efficient compared to the baseline, achieving a 9.1% improvement in macro AUROC and a 34.9% improvement in macro AUPRC using just 10% of the training data. Ablation studies identify key components, such as stochastic depth and preview linear probing, that contribute to the enhanced performance. These findings underscore the potential of post-training strategies to improve ECG foundation models, and we hope this work will contribute to the continued development of foundation models in the ECG domain.
@@
Keywords: Electrocardiogram, Post-training strategy, ECGFounder, Model adaptation, Foundation models
@@ 1 Introduction
Electrocardiography (ECG) is a widely used, non-invasive technique for the screening, diagnosis, monitoring, and risk prediction of cardiovascular diseases. ECG captures the heart’s electrophysiological activity, offering valuable insights into cardiac function and health. In recent years, the rapid evolution of artificial intelligence, particularly deep learning, has revolutionized ECG analysis. Deep neural networks (DNNs) have demonstrated expert-level accuracy in detecting a wide range of cardiac abnormalities, such as arrhythmias (Hannun et al., 2019; Ribeiro et al., 2020). These models are increasingly deployed in clinical settings to assist cardiologists, alleviate diagnostic workloads, and enhance decision-making (Jiang et al., 2024). Furthermore, DNNs have demonstrated the ability to identify subtle signal patterns undetectable by human experts (Somani et al., 2021; Siontis et al., 2021), enabling the diagnosis of conditions like structural heart disease (Poterucha et al., 2025), anemia (Kwon et al., 2020), and diabetes (Kulkarni et al., 2023). This has expanded the diagnostic potential of ECG far beyond its traditional role.
Recently, foundation models have emerged as a transformative approach for building generalizable and transferable models across diverse clinical tasks (Moor et al., 2023). Trained on large-scale datasets, these models can be fine-tuned for various downstream tasks (Bommasani et al., 2021). In the ECG domain, foundation models such as ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ECGFounder (Li et al., 2025) have shown broad applicability across ECG-related tasks. The recently released foundation model, ECGFounder, built on the Harvard-Emory ECG Database (Koscova et al., 2024), which includes over 10 million clinically annotated ECG recordings covering 150 types of cardiac abnormalities, has demonstrated strong performance and generalizability across multiple applications.
@@
Despite significant advancements in ECG foundation models, substantial performance gaps persist when fine-tuned for specific diagnostic tasks, especially compared to recent task-specific models. For example, a recent study (Coppola et al., 2024) reveals that HuBERT-ECG BASE, while performing well across several tasks on the PTB-XL benchmark, falls short when compared to those specialized models (Strodthoff et al., 2021). Encouraged by the open-source efforts of Li et al. (2025), we conducted a series of experiments using the PTB-XL dataset. However, we observed that the fine-tuning performance of even the state-of-the-art foundation model ECGFounder remains suboptimal. This highlights a key issue: when ECG foundation models pre-trained on one large dataset are applied to ECG signals from other data sources, they often fail to match the performance of specialized models. Moreover, compared to recent advanced architectures—such as multi-scale models (Shi et al., 2023) and state-space models (Zhang et al., 2023; Behrouz et al., 2024)—the performance of these foundation models lags significantly, raising concerns about their practical applicability in real-world clinical settings.
@@
To address these limitations, we propose a simple yet effective post-training strategy based on the recently released ECGFounder model. We use the PTB-XL all-label classification benchmark as a performance example. As illustrated in Figure 1, while pre-training on over 7 million ECG recordings brings substantial improvements over training from scratch, the benefits of pre-training are closely matched by our post-training approach, even when applied to randomly initialized models. When applied to the pre-trained ECGFounder, our proposed post-training strategy leads to remarkable performance gains of 3.2% in AUROC and 14.3% in AUPRC (Figures 1(a) and 1(b)). Furthermore, our experiments show that, compared to state-of-the-art task-specific models, our method achieves superior classification performance on several PTB-XL benchmark tasks. These results underscore the complementary advantages of the proposed post-training strategy, demonstrating its ability to address the performance gaps observed in ECG foundation models.
Figure 1: Performance comparison on PTB-XL 71-label classification benchmark. (a) Marco AUROC results comparing different training strategies. (b) Macro AUPRC results under the same experimental setup. “Random Initialization” denotes training from scratch, while “Pre-training” utilizes transferred weights. Both subfigures contrast baseline fine-tuning methods (left) with our proposed post-training approach (right).
@@
To explore the effectiveness of the proposed post-training strategy, we conducted a series of experiments, including varying sample sizes and training length, and an ablation study. By evaluating performance under diverse conditions, we found that the strategy is both stable and sample-efficient, showing its capacity to enhance model performance not only with full datasets but also in scenarios with limited labeled data. These experiments provide valuable insights, demonstrating the promising performance and potential applicability of the method.
@@
    Post-Training Strategy for ECGFounder. We propose a simple yet effective post-training strategy to enhance ECGFounder, a recently released state-of-the-art ECG foundation model. This strategy provides a crucial complement to the original framework, facilitating more efficient adaptation to diverse downstream ECG tasks.
@@
    Significant Performance Gains. Experiments on the PTB-XL benchmark demonstrate that our method outperforms the baseline fine-tuning strategy from the original work across multiple tasks, by 1.2%–3.3% in macro AUROC and 5.3%–20.9% in macro AUPRC. Furthermore, we explore the properties of the proposed method under varying sample sizes and training epochs. Notably, with only 10% of the training data, the proposed method achieves remarkable improvements of 9.1% in macro AUROC and 34.9% in macro AUPRC, highlighting its effectiveness in data-scarce scenarios.
@@
    Superior Performance. Our method surpasses several recent state-of-the-art task-specific and advanced models across multiple ECG classification tasks, highlighting the potential of ECG foundation models when equipped with an effective post-training strategy.
  • 4.
@@ 2.1 ECG Classification and Evaluation Benchmarks
ECG classification is one of the most important tasks in ECG analysis, crucial for many critical applications, including the detection of arrhythmia (Hannun et al., 2019; Ribeiro et al., 2020; Hughes et al., 2021), structural heart disease (Poterucha et al., 2025), anemia (Kwon et al., 2020), and diabetes (Kulkarni et al., 2023). Due to their excellent performance in both accuracy and diagnostic speed (Jiang et al., 2024), deep learning methods have gained significant popularity in these tasks. Successful model architectures in these areas range from convolutional neural networks (CNNs) and Transformer to state-space models and ensemble approaches (Strodthoff et al., 2021; Cheng et al., 2023; Shi et al., 2023; Zhang et al., 2023; Behrouz et al., 2024). Recent advanced architectures have proven effective in modeling the time-series nature of ECG signals, including multi-scale models like MSW-Transformer (Cheng et al., 2023) and MULTIRESNET (Shi et al., 2023), as well as advanced state-space models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024).
However, comparing the performance of various ECG classification methods presents several challenges. While numerous large ECG datasets have been collected, many remain private (Somani et al., 2021; Siontis et al., 2021). Recent efforts to address this limitation have led to the release of several large, publicly available ECG datasets (Clifford, 2024). Nonetheless, a fair comparison requires more than just raw datasets. Clearly defined usage guidelines, such as pre-defined train-test-splits, label sets, and target metrics, are essential to ensure valid comparisons (Strodthoff, 2024). Among these, the PTB-XL dataset (Wagner et al., 2020) stands out as one of the most comprehensive benchmarks, providing essential data for evaluating various ECG classification methods. Furthermore, based on this dataset, Strodthoff et al. (2021) presents benchmark results for several methods (Ismail Fawaz et al., 2020; Wang et al., 2017), providing an important baseline for comparing different ECG classification algorithms.
@@ 2.2 ECG Foundation Models
Recent advances in foundation models for natural language processing have inspired similar approaches for health data (Clifford, 2024). In the ECG domain, several studies have explored building foundation models through large-scale pre-training followed by evaluation across diverse downstream tasks. Representative examples include ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ECGFounder (Li et al., 2025). Specifically, ECG-FM is a Transformer-based model pre-trained on over 1 million 12-lead ECGs from multiple sources; HuBERT-ECG is another Transformer-based model pre-trained on more than 9 million 12-lead ECGs covering 164 cardiovascular conditions; and ECGFounder is a CNN-based model pre-trained on over 7 million 12-lead ECGs annotated with 150 diagnostic categories. Importantly, the codes and pre-trained model weights of all three models have been publicly released.
Comparing the performance of different ECG foundation models would provide valuable insights for practitioners. However, conducting a fair comparison is challenging, especially for ECG foundation models due to their higher computational cost. As discussed in Section 2.1, the PTB-XL benchmark offers a fair platform for evaluating various algorithms. Nonetheless, the pre-training dataset for ECG-FM uses random splits of the PhysioNet 2021 dataset (Reyna et al., 2021; Goldberger et al., 2000), which includes the PTB-XL dataset, potentially limiting its applicability for evaluation on the PTB-XL benchmark. For HuBERT-ECG, the authors have reported performance results on PTB-XL, enabling direct comparison. For ECGFounder, we follow the recommended fine-tuning strategy (Li et al., 2025) based on the released model weights to evaluate performance on the PTB-XL benchmark.
@@ 2.3 Regularization in Deep Learning
Following the general definition of regularization, initialization strategies can be considered a form of regularization. The pre-training paradigm, which has been successful in many ECG tasks (Zhou et al., 2025; Li et al., 2025; Coppola et al., 2024; McKeen et al., 2024), retains the pre-trained weights in the early layers of the model. However, the final classification head is typically initialized randomly. In this paper, we use linear probing to initialize this layer. Furthermore, early stopping can also be regarded as a regularization technique, and it is often used alongside other regularization methods (Goodfellow et al., 2016). We incorporate early stopping in our experiments as well.
@@ 3 Methodology
In this study, we propose a post-training strategy designed to enhance the classification performance of ECGFounder, a recently released state-of-the-art ECG foundation model (Li et al., 2025). The overall framework of the proposed approach is illustrated in Figure 2. Our framework builds upon the publicly available network architecture and pre-trained weights of ECGFounder, and introduces two key stages for downstream classification: an initialization stage and a regularization stage. The design and implementation of each stage are described in detail in the following subsections.
@@
3.1 Overview of ECGFounder
This subsection provides a concise overview of ECGFounder. The model is built upon the CNN-based Net1D architecture (Hong et al., 2020), adapted from RegNet (Radosavovic et al., 2020) and specifically optimized for ECG signal analysis. Its design incorporates stagewise scaling, bottleneck blocks (He et al., 2016), and channel-wise attention (Hu et al., 2018), enabling the model to capture both temporal dynamics and inter-lead spatial dependencies.
For pre-training, ECGFounder was built on part of the Harvard-Emory ECG Database (Koscova et al., 2024), which contains over 10 million ECGs across 150 diagnostic categories, using a multilabel setting that reflects the complexity of real-world clinical practice. To address incomplete annotations, a positive–unlabeled (PU) learning strategy (Zhao et al., 2022) was integrated into the loss function, allowing the model to account for unlabeled but potentially positive samples.
For downstream tasks, ECGFounder supports both linear probing and full fine-tuning, as proposed in Li et al. (2025). These training strategies are optimized to ensure robust performance across diverse ECG applications. Specifically, training is guided by a ReduceLROnPlateau learning rate scheduler with the AdamW optimizer. Model selection is based on validation macro Area Under the Receiver Operating Characteristic Curve (AUROC), aiming to enhance adaptability and reliability.
@@ 3.2 Post-training Strategy
As illustrated in Figure 2, our method builds upon ECGFounder and introduces a two-stage post-training strategy. The strategy is designed to enhance classification performance while requiring minimal modifications to the original model architecture. Specifically, it consists of an initialization stage and a regularization stage, which are described in detail below.
@@ 3.2.2 Regularization Stage
The regularization stage involves full fine-tuning of the entire model following initialization, with multiple regularization techniques incorporated to improve generalization. The core component is the stochastic depth strategy (Huang et al., 2016), which has been shown to effectively reduce redundancy in ECG signals (Zhou et al., 2025). Specifically, stochastic depth is applied to all residual connections in ECGFounder. With a predefined drop-path rate $\lambda$, a binary random mask $r$ is defined as
$r=\mathbb{I}(U<1-\lambda),\quad U\sim\mathcal{U}(0,1),$ (1)
where $\mathcal{U}$ denotes the uniform distribution and $\mathbb{I}(\cdot)$ is the indicator function. Let $\mathbf{X}$ represent the input tensor, $F(\cdot)$ denote the block transformation, and $\widetilde{\mathbf{X}}$ the block output. The stochastic depth formulation during training is expressed as
$\widetilde{\mathbf{X}}=\frac{r}{1-\lambda}\cdot F(\mathbf{X})+\mathbf{X},$ (2)
where $r$ determines whether the transformation $F(\cdot)$ is active $(r=1)$ or skipped $(r=0)$, and the scaling factor $1/(1-\lambda)$ ensures that the expected output is preserved. During inference, the deterministic residual form is used:
$\widetilde{\mathbf{X}}=F(\mathbf{X})+\mathbf{X}.$ (3)
To further stabilize training, Dropout (Hinton et al., 2012; Srivastava et al., 2014) is applied to the activated outputs within all bottleneck blocks. In addition, a cosine annealing learning rate schedule (Loshchilov and Hutter, 2017) and weight decay (Hanson and Pratt, 1988; Loshchilov and Hutter, 2019) are employed to provide smooth decay and enhance robustness during optimization.
3.3 Ease of Implementation
To preserve the simplicity and accessibility of ECGFounder, we deliberately minimize modifications to the original framework. All proposed strategies avoid deep architectural changes and can be readily integrated into the existing implementation. In the initialization stage, our method can be directly applied within the open-source ECGFounder code. In the regularization stage, Dropout, weight decay the cosine annealing learning-rate schedule are standard functions in PyTorch. The stochastic depth operation is also straightforward to implement, requiring only a single-line modification within the bottleneck block:
  • out += identity
is replaced with
  • out = self.drop_path(out) + identity
Here, out and identity are tensors already defined in the implementation of Li et al. (2025), while self.drop_path() represents the stochastic depth strategy employed in this work. Overall, the proposed post-training strategy remains simple yet effective, preserving the ease of use of ECGFounder while providing enhanced performance.

### v2

# Bridging Performance Gaps for ECG Foundation Models: A Post-Training Strategy

@@ Abstract
ECG foundation models are increasingly popular due to their adaptability across various tasks. However, their clinical applicability is often limited by performance gaps compared to task-specific models, even after pre-training on large ECG datasets and fine-tuning on target data. This limitation is likely due to the lack of an effective post-training strategy. In this paper, we propose a simple yet effective post-training approach to enhance ECG foundation models. We evaluate it on a publicly available Transformer-based foundation model. Experiments across multiple ECG tasks show that our method consistently outperforms baseline fine-tuning. On the PTB-XL benchmarks, it improves macro AUROC by 0.7%-8.9% and macro AUPRC by 23.3%-77.9%, also outperforming several recent state-of-the-art approaches, including task-specific and advanced architectures. Further analyses demonstrate improved training dynamics and data efficiency, with only 30% of the training data outperforming the baseline trained on the full dataset. Ablation studies highlight the importance of stochastic depth and preview linear probing. These findings underscore the potential of post-training strategies to improve ECG foundation models, and we hope this work will contribute to the continued development of foundation models in the ECG domain.
@@
Keywords: Electrocardiogram, Post-training strategy, Model adaptation, Foundation models
@@ 1 Introduction
Foundation models have recently emerged as a transformative paradigm in medical artificial intelligence, enabling the development of generalizable and transferable models across diverse clinical tasks (Moor et al., 2023; Bommasani et al., 2021). Leveraging large-scale pretraining, these models can be efficiently adapted to various downstream applications with limited task-specific data, showing great promise in biomedical analysis. In the electrocardiogram (ECG) domain, several foundation models, such as ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ST-MEM (Na et al., 2024), have been introduced to advance ECG-based tasks. Trained on extensive ECG datasets, these models aim to capture universal cardiac representations and support multiple downstream applications, such as arrhythmia classification (Hannun et al., 2019; Ribeiro et al., 2020; Jiang et al., 2024) and disease risk prediction (Somani et al., 2021; Siontis et al., 2021; Poterucha et al., 2025; Kwon et al., 2020; Kulkarni et al., 2023).
@@
Despite their promise, ECG foundation models still face notable limitations. When fine-tuned for specific diagnostic tasks, their performance often lags behind task-specific or advanced architectures such as multi-scale (Shi et al., 2023) and state-space models (Zhang et al., 2023; Behrouz et al., 2024). For instance, HuBERT-ECG-BASE performs competitively across several PTB-XL tasks (Strodthoff et al., 2021) yet remains inferior to many specialized models (Coppola et al., 2024). Benefiting from the recent open-source releases of many ECG foundation models, we were able to conduct a series of experiments on the PTB-XL dataset. Our results revealed similar findings: although ECG foundation models demonstrate broad adaptability across diverse tasks, their fine-tuning performance remains suboptimal when transferred to datasets from different sources, highlighting concerns about their practical applicability in real-world clinical settings.
@@
To address these limitations, we propose a simple yet effective post-training strategy to enhance ECG foundation models. Using the performance of a Transformer-based model (Na et al., 2024) on the PTB-XL all-label classification benchmark as an example, we demonstrate that our strategy yields remarkable performance gains of 5.2% in macro AUROC and 34.9% in macro AUPRC. Furthermore, compared with recent state-of-the-art task-specific and advanced architectures, our approach demonstrates superior classification performance across multiple tasks. These results underscore the complementary advantages of the proposed post-training strategy, demonstrating its potential to address the performance gaps observed in ECG foundation models.
@@
    Post-Training Strategy for General ECG foundation models. We propose a simple yet effective post-training strategy to enhance ECG foundation models. This strategy serves as a crucial complement to existing fine-tuning frameworks, enabling more efficient and robust adaptation across diverse downstream ECG classification tasks.
@@
    Significant Performance Gains. Experiments demonstrate that our approach consistently outperforms the baseline fine-tuning strategy across multiple datasets. On the PTB-XL benchmarks, our method achieves improvements of 0.7%-8.9% in macro AUROC and 23.3%-77.9% in macro AUPRC, also outperforming many recent state-of-the-art approaches, including task-specific and advanced architectures.
@@ 2.1 ECG Classification and Evaluation Benchmarks
ECG classification is one of the most important tasks in ECG analysis, crucial for many critical applications, including the detection of arrhythmia (Hannun et al., 2019; Ribeiro et al., 2020; Hughes et al., 2021), structural heart disease (Poterucha et al., 2025), anemia (Kwon et al., 2020), and diabetes (Kulkarni et al., 2023). Due to their excellent performance in both accuracy and diagnostic speed (Jiang et al., 2024), deep learning methods have gained significant popularity in these tasks. Successful model architectures in these areas range from convolutional neural networks (CNNs) and Transformer to state-space models and ensemble approaches (Strodthoff et al., 2021; Cheng et al., 2023; Shi et al., 2023; Zhang et al., 2023; Behrouz et al., 2024). The task-specific and advanced architectures have proven effective in modeling the time-series nature of ECG signals, including the residual models like resnet1d101 (Strodthoff et al., 2021), resnet1d_wang (Wang et al., 2017), the multi-scale models like MSW-Transformer (Cheng et al., 2023) and MULTIRESNET (Shi et al., 2023), as well as advanced state-space models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024).
Despite rapid methodological progress, fair comparison across ECG classification methods remains challenging due to the lack of benchmark datasets. Although numerous large-scale ECG datasets have been collected worldwide, many are not publicly available (Somani et al., 2021; Siontis et al., 2021). To address this limitation, several large open-access ECG datasets have been released in recent years (Clifford, 2024), including PTB-XL (Wagner et al., 2020), Chapman-Shaoxing (Zheng et al., 2020b), and Ningbo (Zheng et al., 2020a). Among these, PTB-XL has become the one of the most widely adopted benchmarks due to its comprehensive usage guidelines, including predefined train-test splits, standardized label sets, and recommended evaluation metrics (Strodthoff, 2024). These resources have established PTB-XL as a widely acceptable benchmark for ECG classification, enabling reproducible and comparable evaluation across diverse studies (Strodthoff, 2024; Strodthoff et al., 2021).
@@ 2.2 ECG Foundation Models
Recent advances in foundation models for natural language processing have inspired similar approaches for health data (Clifford, 2024). In the ECG domain, several studies have explored building foundation models through large-scale pre-training followed by evaluation across diverse downstream tasks. Recent examples include ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ECGFounder (Li et al., 2025). Noting that the predominant architecture of existing foundation models is Transformer-based, this paper focuses on Transformer-based models, although the proposed strategy may also be applicable to CNN-based models, such as ECGFounder. Specifically, ECG-FM is a Transformer-based model pre-trained on over 1 million 12-lead ECGs from multiple sources; HuBERT-ECG is another Transformer-based model pre-trained on more than 9 million 12-lead ECGs covering 164 cardiovascular conditions. Another line of ECG foundation models has emerged from research on self-supervised learning for ECG signals. For instance, Na et al. (2024) proposed ST-MEM, a Transformer-based SSL framework that explicitly captures the spatiotemporal dependencies inherent in ECG signals. ST-MEM was pre-trained on 188,480 ECGs and evaluated across various downstream tasks. Notably, all four models have publicly released their source code and pre-trained weights, promoting transparency and reproducibility in ECG foundation model research.
Comparing the performance of different ECG foundation models would provide valuable insights for practitioners. However, conducting a fair comparison is challenging, especially for ECG foundation models due to their higher computational cost. As discussed in Section 2.1, the PTB-XL benchmark offers a fair platform for evaluating various algorithms. Nonetheless, the pre-training dataset for ECG-FM uses random splits of the PhysioNet Challenge 2021 dataset (Reyna et al., 2021; Goldberger et al., 2000), which includes the PTB-XL dataset, potentially limiting its applicability for evaluation on the PTB-XL benchmark. By contrast, HuBERT-ECG reports benchmark results directly on PTB-XL, enabling more straightforward comparisons. In this work, we focus on developing a general post-training strategy for ECG foundation models. Given that Transformer-based architectures are widely adopted in contemporary foundation models, we select ST-MEM (Na et al., 2024) as a representative example. Throughout this paper, we denote such Transformer-based foundation models as Transformer-FM, emphasizing that the proposed post-training strategy is general and not specifically tailored to any single model.
@@ 2.3 Regularization in Deep Learning
Following the general definition of regularization, initialization strategies can be considered a form of regularization. The pre-training paradigm, which has been successful in many ECG tasks (Na et al., 2024; Coppola et al., 2024; McKeen et al., 2024), retains the pre-trained weights in the early layers of the model. However, the final classification head is typically initialized randomly. In this paper, we use linear probing to initialize this layer. Furthermore, early stopping can also be regarded as a regularization technique, and it is often used alongside other regularization methods (Goodfellow et al., 2016). We incorporate early stopping in our experiments as well.
@@ 3 Methodology
In this study, we propose a simple yet effective post-training strategy to enhance the classification performance of ECG foundation models. The proposed strategy is model-agnostic, and a Transformer-based foundation model is selected as a representative example to demonstrate its effectiveness. Notably, the components of the proposed strategy are easy to implement and can be readily integrated into a wide range of ECG models.
The overall framework of the proposed approach is illustrated in Figure 1. It builds upon pre-trained ECG foundation models, and introduces two key stages for downstream classification: an initialization stage and a regularization stage. The design and implementation of each stage are described in detail in the following subsections.
@@
3.1 Overview of the Transformer-based example
Transformer-FM (Na et al., 2024) is a Transformer-based foundation model adapted from the Masked Autoencoder (MAE) framework (He et al., 2022), specifically designed to capture the spatio-temporal relationships in multi-lead ECG signals. It segments ECG sequences along both the temporal and lead dimensions, employing a lead-wise shared decoder and lead-specific embeddings to effectively model inter-lead dependencies. For pre-training, Transformer-FM was trained on a hybrid dataset containing 188,480 ECGs from the Chapman-Shaoxing (Zheng et al., 2020b), Ningbo (Zheng et al., 2020a), and CODE-15 (Ribeiro et al., 2021) datasets. For downstream tasks, it supports both linear probing and full fine-tuning, following the MAE training protocol with a cosine learning rate decay and the AdamW optimizer.
@@ 3.2 Post-training Strategy
As illustrated in Figure 1, our method builds upon pre-trained ECG foundation models and introduces a two-stage post-training strategy. The strategy is designed to enhance classification performance while requiring minimal modifications to the original model architecture. Specifically, it consists of an initialization stage and a regularization stage, which are described in detail below.
@@ 3.2.2 Regularization Stage
The regularization stage involves full fine-tuning of the entire model following initialization, with multiple regularization techniques incorporated to improve generalization. The core component is the stochastic depth strategy (Huang et al., 2016), which has been shown to effectively reduce redundancy in ECG signals (Zhou et al., 2025). Specifically, stochastic depth is applied to all residual connections in Transformer-FM. To further stabilize training, Dropout (Hinton et al., 2012; Srivastava et al., 2014) is applied to the feed-forward linear layers for Transformer-FM. In addition, a cosine annealing learning rate schedule (Loshchilov and Hutter, 2017) and weight decay regularization (Hanson and Pratt, 1988; Loshchilov and Hutter, 2019) are employed to ensure smooth learning rate decay and enhance optimization robustness for Transformer-FM.